A gaseous mixture of h2 and O2 at 60°C and total pressure of 2.5 atmosphere contain 75 mole % h2 what are the partial pressure of both gases in the mixture?


1
Expert's answer
2021-02-22T06:09:12-0500

If the mole fraction of H2 is 75% (or 0.75), the mole fraction of O2 equals 1 - 0.75 = 0.25. Therefore,

"P_{H_2}=\chi_{H_2}\times{P_{total}}=0.75\times2.5=1.9atm"

"P_{O_2}=\chi_{O_2}\times{P_{total}}=0.25\times2.5=0.63atm"


Answer: H2 - 1.9 atm; O2 - 0.63 atm